X + f = 120
x + p = 230
f X 6 = p

Answer by richwmiller(17219)

f X 6 = p
X + f = 120
Don't use X for multiplication and as a variable
X + f = 120
x + p = 230
X and x are two different variables.
6f = p
x + p = 230
x + f = 120
subtract
p-f=110
6f=p
6f-f=110
5f=110
f=22
x=98
p-f=110
p-22=110
p=132